OpenStudy (anonymous):

solve: 4/n - 1/2 = -6/(3n)

OpenStudy (mathmale):

I'd first identify the LCD (lowest common denominator) of this equation involving fractions. Then I'd multiply every term by this LCD. The resulting equation would then have no fractions and thus be easier to solve. Solve this equation.

OpenStudy (mathmale):

Your LCD is ... ?

OpenStudy (anonymous):

I thought of that too, but how can i find this LCD?

OpenStudy (anonymous):

oh i see, if i find the lcm and simplify 6/3n i get 2/n. Then i can find a lcd

OpenStudy (anonymous):

how? N is a factor, but i 2 and 3? "6"?

OpenStudy (mathmale):

Not sure you need an "lcm." Look at the denominators of the several fractions in your equation. They are n, 2 and 3n, or, in factored form, n, 2 and (3)(n). As you say, n is one factor of the LCD. 2 is also, and 3 is also. Simply multiply these three together, and then you've got your LCD.

OpenStudy (anonymous):

alright, ill try it.

OpenStudy (mathmale):

(n)(2)(3) = ? This is your LCD. Now multiply each term of your original equation by this LCD. Reduce the resulting equation. It will no longer have fractions. Solve this new equation.

OpenStudy (anonymous):

i think i did this wrong... 17-n=12? (3)(2)(n)*4/n - 1/2*(3)(2)(n) = 6/3n*(3)(2)(n)? 12+8 - 3 - n = 12

OpenStudy (mathmale):

Why? If 17-n=12, add n to both sides to force n to be positive, and then subtract 12 from both sides. Result? n= ?? Now test this n value by substituting it back into the orginal equation. Is the eq'n now true or false?

OpenStudy (anonymous):

false. it doesn't work. n must equal to 12.

OpenStudy (mathmale):

Hold a moment. It's most likely an arithmetic error.

OpenStudy (mathmale):

I've done the problem myself and find that n does indeed equal 12. What could I do at this point to make this problem solution clearer for you?

OpenStudy (mathmale):

\[6n \left( 4/n - 1/2 =-6/(3n) \right)=>?\]

OpenStudy (mathmale):

I get \[24-3n=-12.\]